A fish was caught whose tail weighed 9 pounds. His head weighed as much as his tail and half his body,and his body weighed as much as his head and tail.How much did the fish weigh?
Answer by flame8855(424)

t is weight for tail
h is the weight for head
b is the weight for body
t =9
h = t + b/2
h = 9 +b/2
b = h+t = h + 9
h = 9 + h/2 + 9/2
h = 27
b = 36